Prove that... sin0-cos0+1/sin0+cos0-1= 1/sec0-tan0 … plz prove it and send the answer....thnx

Grade:10

1 Answers

saurav kumar
6 years ago
LHS=sin0-cos0+1/sin0+cos0-1
on dividing by cos0 up and down,then
tan0+se0-1/tan0-sec0+1
= sec0+tan0-(sec20-tan20)/tan0-sec0+1
= sec0+tan0-{(sec0+tan0)(sec0-tan0)}/tan0-sec0+1
=sec0+tan0(1-sec0+tan0)/tan0-sec0+1
=sec0+tan0*(sec0-tan0)/(sec0-tan0)
=(sec20-tan20)/(sec0-tan0)
=1/sec0-tan0)
=RHS